OpenStudy (anonymous):

If you have lifted 10 pounds 2 feet, you have done___ foot-pounds of work.

OpenStudy (anonymous):

20, Amiright?

OpenStudy (anonymous):

Yup. Technically, this is physics, but yup.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The equation for work is W = F * D, where F is force in lbs and D is distance is feet. F = 10 * 2 = 20. So you are correct.